OpenStudy (anonymous):

How many moles of zinc are in 11 g?

OpenStudy (astrophysics):

m = n * M, where M is the molar mass

OpenStudy (astrophysics):

so the molar mass of zinc is 65/38 g/mol so \[m = n \times M \implies n = \frac{ m }{ M }\]

OpenStudy (butterflydreamer):